OpenStudy (megannicole51):

A pyramid of constant density gm/cm^3 has a square base of side length 40cm, and a height of 10cm. What is the mass of the pyramid and center mass.

OpenStudy (anonymous):

what is the volume of the pyramid ? after you calculate the volume V the mass M will be : M = V * density

OpenStudy (megannicole51):

v=(1/3)*b*h v=(1/3)*10*40 v=400/3

OpenStudy (megannicole51):

mass=(400/3)*density and idk what the density is

OpenStudy (anonymous):

i think they said : 1 gm/cm^3

OpenStudy (megannicole51):

there isnt a 1 written in the question

OpenStudy (megannicole51):

it just has the density symbol in front of that

OpenStudy (anonymous):

so you can give numerical answer.. it will be dependent on the density value

OpenStudy (megannicole51):

not sure what u mean...

OpenStudy (anonymous):

if the density is x then M = (400x)/3

OpenStudy (megannicole51):

theres the answer....idk how they got it

OpenStudy (anonymous):

v=(1/3)*b*h b = 40 * 40 (base area) h = 10

OpenStudy (anonymous):

and as you can see the M is this V times the density symbol

OpenStudy (megannicole51):

yeah my professor made this so much more complicated...so then how do we find the center of mass?

OpenStudy (anonymous):

so for the center of mass we need the intergal. first you need to understand that the x and y of the center of mass is 0 (due to symmetry)

OpenStudy (anonymous):

for the z point you need the integral

OpenStudy (megannicole51):

a=0 b=10

OpenStudy (anonymous):

oh what i said is written there already

OpenStudy (megannicole51):

and isnt something (1/M)?

OpenStudy (anonymous):

yes

OpenStudy (megannicole51):

oksy keep going:)

OpenStudy (anonymous):

ok so now you need triple integral zdxdydz

OpenStudy (anonymous):

do you know what are the limits of integration ?

OpenStudy (megannicole51):

yes

OpenStudy (anonymous):

good - you better write y and x limits as functions of z

OpenStudy (megannicole51):

zdxdydz not sure what or why u wrote this

OpenStudy (anonymous):

triple integral

OpenStudy (megannicole51):

whats the formula to find the center of the mass? and why do we need a triple integral?

OpenStudy (anonymous):